# Using $k$-medoids for distributed approximate similarity search with arbitrary distances

### v1

@@ Abstract
This paper presents GMASK, a general algorithm for distributed approximate similarity search that accepts any arbitrary distance function. GMASK requires a clustering algorithm that induces Voronoi regions in a dataset and returns a representative element for each region. Then, it creates a multilevel indexing structure suitable for large datasets with high dimensionality and sparsity, usually stored in distributed systems. Many similarity search algorithms rely on $k$-means, typically associated with the Euclidean distance, which is inappropriate for specific problems. Instead, in this work we implement GMASK using $k$-medoids to make it compatible with any distance and a wider range of problems. Experimental results verify the applicability of this method with real datasets, improving the performance of alternative algorithms for approximate similarity search. In addition, results confirm existing intuitions regarding the advantages of using certain instances of the Minkowski distance in high-dimensional datasets.
Keywords Approximate similarity search  $\cdot$ k-medoids  $\cdot$ arbitrary distance function
@@ 1 Introduction
The goal of similarity search is to find elements from a dataset, represented by feature vectors in a multidimensional space, similar to a given query object according to a particular similarity or dissimilarity measure. Numerous data indexing methods, collectively known as access methods, have been proposed to speed up this task. Nevertheless, the results remain strongly conditioned by the problems inherent to the representation of datasets in high-dimensionality and high-sparsity spaces. These algorithms are frequently conceived to work with data having spatial components or represented in a vector space with few dimensions. However, they become unusable in high-dimensional spaces, where many features characterise each element due to the Curse of Dimensionality [1]. As a result, the distance between elements tends to become increasingly uniform, making it challenging to differentiate between them based solely on their distances or when the number of dimensions significantly exceeds the quantity of data points. Moreover, several studies indicate that in datasets with over 20 dimensions, exact similarity search methods rarely outperform exhaustive scanning.
Owing to this constraint, novel approaches have emerged in recent years to swiftly identify enough related items within a reasonable time frame compared to linear scanning, often referred to as approximate similarity search. One of the newest approaches is MASK (Multilevel Approximate Similarity search with k-means) [2], a novel indexing method that involves an unconventional application of the $k$-means partitioning algorithm [3, 4] to create a multilevel index structure for approximate similarity search.
Traditionally, the input to $k$-means is always a fixed number of $k$ prototype points or centroids. Then, following an iterative process, the dataset is divided into $k$ groups (Voronoi cells), assigning each point to its closest centroid according to a given distance function, so that each centroid represents all of the elements of that cluster. In contrast, leveraging current powerful computing capabilities, $k$-means can be forced to generate a high number of centroids that inundate the dataset. Then, these centroids become a new dataset to be grouped at the upper level, repeating this process until the complete construction of the search tree is achieved. This bottom-up creation of the multilevel index enables the distribution of this task among different computing nodes, which also store distinct data partitions.
Although it employs an innovative approach that capitalises on the benefits offered by contemporary computing systems and distributed platforms, the performance of this indexing method remains significantly impacted by the inherent constraints of the $k$-means algorithm. The choice of the distance function significantly impacts the outcome of $k$-means clustering, as it determines how the similarity between data points is calculated and can lead to different outcomes in terms of the shape and separation of the clusters.
Until now, experiments conducted with $k$-means have relied on the Euclidean distance to calculate relative distances between points, following the conventional inclination to associate the concept of distance with that used in geospatial contexts. Nonetheless, relying on the Euclidean distance as the default option in indexing structures and algorithms when working with high-dimensional data, despite its widespread use, can lead to sub-optimal results. This is because, in high-dimensional spaces, the discriminating power of the $L_{p}$ norm decreases for higher values of $p$ due to their specific features (such as dispersion, outliers, and concentration of points around certain values), leading to poor differentiation between proximate and distant elements.
@@
This paper introduces GMASK (General Multilevel Approximate Similarity Search with $k$-medoids), a novel algorithm that facilitates the incorporation of arbitrary distances by employing data partitioning algorithms that induce Voronoi regions in a dataset and yield a representative element, exemplified by $k$-medoids. The simple introduction of $k$-medoids turns GMASK into a versatile algorithm capable of adapting to the characteristics of any dataset (high dispersion, high dimensionality, outliers, etc.) and seek for the optimal results. This approach allows using multiple $\ell_{p}$ norms or even other distance functions, making it compatible with a broader spectrum of problems. To comprehensively evaluate its performance, an extensive series of experiments will be conducted across datasets varying in type, size, dimensionality, and sparsity.
@@
This paper adheres to the following structure. Section “Background and related work” reviews previous research related to similarity search. Section “Methodology” describes GMASK, emphasising the potential for enhancing the algorithm’s performance by exploring the application of other distance functions beyond the Euclidean one, like other Minkowski distances. The findings of four empirical experiments carried out over real-world datasets with quite diverse attributes are presented in Section “Experimental results”. These tests compare GMASK performance with other well-known algorithms, such as FLANN and PyNNdescent. Finally, Section “Conclusions and future work” recaps the main conclusions of this work and suggests promising lines for future work.
@@ 1.1 Contributions
The potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a wider range of dissimilarity measures [5], justifies our efforts to explore alternative Minkowski distances for data indexing. This endeavour is aimed at challenging the misconception we have identified and fostering a deeper understanding of the significant impact of using alternative distance metrics when working with high-dimensional data.
  • GMASK allows the use of different clustering algorithms for the bottom-top construction of the index tree, such as $k$-medoids, a broader form of the $k$-means algorithm that offers parameterisation with several distance metrics, in particular Minkowski distances.
  • Evaluating the performance of different $L_{p}$ norms or Minkowski distances for approximate $k$-nearest neighbour search. This assessment involves testing GMASK against other state-of-the-art methods to illustrate that the selection of a suitable distance function, while considering the features of the target dataset, has a significant impact on performance when searching for reasonably similar regions.
@@ 2.1 Similarity search
Similarity search, often referred to as proximity search [6, 7, 8], is a cornerstone for applications in many different fields, such as databases [9], information retrieval [10], computer vision [11] and bioinformatics [12, 13], among others. It aims to find elements that are similar (close) to a given query object, according to some measure of similarity or, conversely, dissimilarity.
Thus, given a query object $q\in\mathbb{X}$, the similarity search tackles some of the following scenarios, represented in Figure 1 [7]:
  • Point query: Finding elements within $\mathbb{X}$ that have exactly the same feature values as $q$.
  • Range query: Identifying a subset of elements $\{o_{i}\}\in\mathbb{X}$ whose feature values do not exceed a given similarity threshold $r$ around $q$ such that $\forall\{o_{i}\},\delta(o_{i},q)\leq r$.
  • Nearest-neighbour query: Finding elements in $\mathbb{X}$ whose features are the closest to the feature vector representing $q$. In this case, the goal might be to retrieve the single element (set of $k$ elements) closest to $q$, known as the nearest neighbour ($k$-nearest neighbours) and denoted by $NN_{q}$ ($k-NN_{q}$). Identifying the $k$-nearest neighbours of a given object is one of the most widely used classification methods in machine learning.
@@
Many techniques for data indexing, collectively known as access methods, can speed up these queries. These methods create an index structure that partitions the domain $\mathbb{X}$ into different regions based on a distance function. Then, we can leverage these regions to solve similarity queries. There is a wide range of access methods, including exact approaches, approximate solutions and variants tailored for distributed computing systems.
@@ 2.2 Algorithms for similarity search
Exact similarity search methods identify the completely accurate result set for a query by using the absolute positions of objects to group elements and search for similar instances. These methods are categorised as Point Access Methods (PAM) for point-based elements and Spatial Access Methods (SAM) for elements with spatial dimensions, such as lines, polygons or higher-dimensional polyhedra. The general acronym SAM is commonly used in the research literature to refer to both categories of indexing strategies in vector spaces [10]. These approaches are most effective with elements represented in a relatively small number of dimensions, making Curse of Dimensionality one of the biggest challenges in data science today.
@@
Metric Access Methods (MAM) provide a general framework for similarity search problems within metric spaces [6, 8], exploiting the triangle inequality to partition the metric space into subspaces, effectively filtering out portions of the dataset that cannot contain valid results. On the one hand, MAM exact search can be expensive in terms of computational time and updating the index structure due to dynamic data. On the other hand, MAM indexes have serious problems in efficiently reducing the search for candidate elements in a query when dealing with high-dimensional data. There are many MAM for queries in metric spaces, some of which use clustering algorithms, such as $k$-means [4] or $k$-medoids [14], to recursively partition the entire set of elements creating a hierarchical structure (tree) of nested clusters.
Approximate indexing methods are solutions that strike a balance between accuracy and speed, making them suitable for both vector and metric spaces [15]. Although these methods may introduce some errors into the results, they also provide a faster response time by calculating fewer distances. As a result, these methods prove feasible in scenarios where exact methods cannot provide a fast response, such as in high-dimensional problems or large datasets, where exact methods default to a full scan. Some of the most widely used exact algorithms are compared and summarised in [16], which also introduces a comprehensive taxonomy to classify them according to relevant characteristics.
Rather than computing the exact result set for a query, approximate similarity search algorithms adopt various strategies (e.g., modifying the representation space, reducing the number of pair distances to compute) to identify a reduced set of candidate items that are likely to be the most similar. FastMap, introduced in [17], is one of the first algorithms to be developed for approximate similarity search in metric spaces. It is based on a method to obtain a fast projection of elements in the original dataset into a user-defined $k$-dimensional space, preserving the distances between pairs of elements in order to achieve an approximate response through dimensionality reduction. In this category of algorithms, MASK [2] emerged as a novel indexing approach characterised by a multilevel design tailored for approximate similarity search. It features an innovative utilisation of the $k$-means partitioning algorithm to quickly reduce the number of checked regions when searching for results.
@@ 2.3 Using different distance functions
The dissimilarity value captures the notion of closeness between two objects, which is interpreted as a distance in a suitable space. A distance function that avoids introducing any correlation between features, such as the Minkowski distance [18] or $\ell_{p}$ norm, can quantify this value [17]. The Manhattan or city block distance ($\ell_{1}$ norm), the Euclidean distance ($\ell_{2}$ norm), and the Chebyshev distance ($\ell_{\infty}$ norm) are typical instances of this family. As many multidimensional indexing structures were initially proposed in the context of spatial systems, many indexing structures and algorithms have commonly relied on the Euclidean distance, derived from its conventional use in two- or three-dimensional spatial applications.
@@
However, several negative consequences and challenges arise when dealing with high dimensional problems, an effect widely known as the Curse of Dimensionality. As the number of dimensions increases, the volume of the space grows exponentially, and the density of points in the space decreases. This sparsity poses a challenge in accurately measuring distances and similarities between data points, since proximity becomes less meaningful. Consequently, points can be very far apart, even though they are close according to a metric such as Euclidean distance. This effect increases the computational complexity of algorithms and operations performed in high-dimensional spaces, often resulting in slower and less efficient computations. For this reason, several authors have pointed out that, due to the Curse of Dimensionality, it is very costly to find the exact nearest neighbour in high-dimensional Euclidean spaces [19, 20].
@@
As a result, although Euclidean spaces reflect our natural conception of space, this natural assumption is challenging to extrapolate when dealing with high-dimensional data. Moreover, traditional clustering algorithms, such as $k$-means, often specify the Euclidean distance as the only appropriate metric without considering other possibilities. This choice brings in a self-imposed limitation that may produce results far from the best [21, 22, 5].
@@
In recent years, several surveys [23, 24, 25] have addressed the evaluation of various metrics for similarity searches across spaces of diverse topology, a burgeoning topic offering insights into the most suitable similarity measures for different problems and contributing to the ongoing development of new methods for similarity search.
@@
Therefore, we aim to introduce GMASK, an algorithm that, being based on a generalisation of earlier concepts previously introduced by MASK, signifies a substantial departure from the latter. GMASK enables the integration of arbitrary distances through the use of data partitioning algorithms that generate Voronoi regions and assign a prototype for each of these regions. Specifically, the utilisation of the $k$-medoids algorithm allows for the incorporation of multiple $\ell_{p}$ norms, facilitating parametrisation with different distance metrics and the execution of validation experiments. Our findings demonstrate the feasibility of employing diverse distances beyond the Euclidean metric, highlighting the potential for achieving superior results in data indexing, contingent upon the specific characteristics of the dataset.

### v2

# A general framework for distributed approximate similarity search with arbitrary distances

@@ Abstract
Similarity search is a central problem in domains such as information management and retrieval or data analysis. Many similarity search algorithms are designed or specifically adapted to metric distances. Thus, they are unsuitable for alternatives like the cosine distance, which has become quite common, for example, with embeddings and in text mining. This paper presents GDASC (General Distributed Approximate Similarity search with Clustering), a general framework for distributed approximate similarity search that accepts arbitrary distances. This framework can build a multilevel index structure, by selecting a clustering algorithm, the number of prototypes in each cluster and any arbitrary distance function. As a result, this framework effectively overcomes the limitation of using metric distances and can address situations involving cosine similarity or other non-standard similarity measures. Experimental results using $k$-medoids clustering in GDASC with real datasets confirm the applicability of this approach for approximate similarity search, improving the performance of extant algorithms for this purpose.
Keywords Data indexing  $\cdot$ Approximate similarity search  $\cdot$ Arbitrary distance  $\cdot$ Clustering  $\cdot$ Unsupervised learning
@@ 1 Introduction
The goal of similarity search is to find elements from a dataset, represented by feature vectors in a multidimensional space, similar to a given query object according to a similarity or dissimilarity measure. Numerous data indexing methods, collectively known as access methods, have been proposed to speed up this task. Nevertheless, their performance remains strongly conditioned by complications inherent to dataset representation in high-dimensional and sparse spaces. These algorithms are often designed to handle data with spatial components or to operate within a low-dimensional vector space. However, they become unusable in high-dimensional spaces, where each element is characterised by numerous features, due to the Curse of Dimensionality Böhm et al. [1]. In such spaces, the distance between elements becomes increasingly uniform, making it challenging to differentiate between them. Analogous issues arise when the number of dimensions significantly exceeds the number of data elements. Moreover, several studies [2] indicate that in spaces exceeding 10 dimensions exact similarity search methods rarely outperform exhaustive scanning.
To address these challenges, researchers have explored various techniques to reduce the dimensionality or transform the feature space to make high-dimensional data easier to manage. A common approach is using dimensionality reduction techniques such as Principal Component Analysis (PCA) and t-Distributed Stochastic Neighbour Embedding (t-SNE) [3]. These techniques project data into lower-dimensional spaces while preserving as much variance or similarity structure as possible. However, these methods are not sufficient, especially if the intrinsic dimensionality remain high after the conversion.
At the same time, approximate similarity search methods have become increasingly important since they attain a practical balance between accuracy and computational efficiency. Methods such as Locality-Sensitive Hashing (LSH) [4] and Random Projection Trees [5] aim to quickly identify nearest neighbours by mapping high-dimensional data into a space where proximity is preserved with high probability. These methods have proven effective in reducing the computational burden associated with high-dimensional similarity searches, although they introduce some trade-offs in terms of accuracy.
Furthermore, hierarchical and partitioning-based methods have also been proposed for multidimensional data search. Methods such as the Ball-Tree [6] or the KD-Tree [7] divide the data space into hierarchical structures, accelerating queries by reducing the number of required distance calculations. However, the performance of these structures may degrade as dimensionality increases and, hence, more robust indexing structures must be used for high-dimensional spaces.
Recent advances have focused on using machine learning models to improve similarity search in high-dimensional spaces. One such approach is deep learning embeddings, where neural networks are trained to map high-dimensional data into lower-dimensional vectors that preserve similarity relationships, which have offered significant improvements in different applications [8]. Although effective, these methods also require significant computing resources and extensive training data to achieve optimal performance.
Finally, the development of similarity search methods highlights a permanent trade-off between computational efficiency and accuracy in addressing high-dimensional data. Future research is likely to continue focusing on hybrid approaches that combine traditional indexing methods with advanced machine learning techniques to push the boundaries of what can be achieved in high-dimensional similarity search. In this regard, it is also worth noticing the emergence of new datasets demanding non-metric distances. In practice, this additional prerequisite discards many existing methods for exact or approximate similarity search that has been conceived specifically for metric spaces.
@@
To overcome these limitations, this paper introduces GDASC, a novel framework that allows arbitrary distances by employing different clustering algorithms that can assign a representative element to each cluster. These properties turn GDASC into a versatile tool that can be adapted to the specific traits of any dataset (high dispersion, high dimensionality, outliers, etc.), seeking for optimal results. This approach can utilise multiple $\ell_{p}$ norms besides the Euclidean distance or even other distance functions, such as cosine distance, making it suitable for a broader spectrum of problems. To comprehensively evaluate its performance, an extensive series of experiments will be conducted across datasets varying in type, size, dimensionality, and sparsity.
@@
The paper adheres to the following structure. Section 2 reviews previous research related to similarity search. Section 3 describes GDASC, emphasising the potential for enhancing the algorithm’s performance by exploring the application of other distance functions beyond the Euclidean one. The findings of four empirical experiments carried out over real-world datasets with quite diverse attributes are presented in Section 4. These tests compare GDASC performance with other well-known algorithms, such as FLANN and PyNN. The implications of using arbitrary distances and different clustering algorithms in similarity searches are discussed in Section 5. Finally, Section 6 recaps the main conclusions of this work and suggests promising lines for future work.
@@ 1.1 Contributions
The potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a wider range of dissimilarity measures [9], justifies our efforts to explore alternative distance metrics for data indexing. This endeavour is aimed at challenging the misconception we have identified and fostering a deeper understanding of the significant impact of using alternative distance metrics when working with high-dimensional data [10].
  • Proposing GDASC, a versatile framework for data indexing that supports the use of arbitrary distances by employing different clustering algorithms for the bottom-up construction of the index tree, such as $k$-medoids.
  • Evaluating the performance of multiple $\ell_{p}$ norms or Minkowski distances as well as cosine distance for approximate $k$-nearest neighbour search. This assessment involves testing GDASC against other state-of-the-art methods to illustrate that the selection of a suitable distance function, while considering the features of the target dataset, significantly impacts performance when searching for reasonably similar regions.
@@ 2.1 Similarity search
Similarity search [11, 12, 13] is an essential technique underpinning applications in a wide range of fields, including but not limited to information retrieval [14], computer vision [15], machine learning [16] and bioinformatics [17, 18]. It is an information retrieval technique involving a set of objects characterised by feature vectors and represented as points in a multidimensional space. The goal is to find the closest (most similar) object to a given query, typically determined by some similarity or dissimilarity measure, often expressed as a distance between two objects.
Let $\mathcal{X}$ denote a set of objects represented by feature vectors in a multidimensional space and $\delta:\mathcal{X}\times\mathcal{X}\mapsto\mathbb{R}$ is a function that assigns a real number $\delta(a,b)$ to each pair of objects $a,b\in\mathcal{X}$, indicating the distance between them. Thus, given a query object $q\in\mathbb{X}$, similarity search can deal with different scenarios, as depicted in Figure 1 [12]:
  • Range query: Finding all objects of $\mathbb{X}$ within a specified threshold around $q$. Formally, retrieving $S=\{o_{i}\mid\delta(o_{i},q)\leq r,\forall o_{i}\in\mathbb{X}\}$.
  • Nearest-neighbour query: Finding the set of objects in $\mathbb{X}$ that are closest to $q$, that is, $S=\{o_{i}\in\mathbb{X}\mid\delta(o_{i},q)\leq\delta(o_{n},q),\forall o_{n}\in\mathbb{X}\}$. The $k$-nearest neighbours method is recognised as the primary approach in this type of search, involving the identification of the $k$ closest objects to $q$ within $\mathbb{X}$. Identifying the $k$-nearest neighbours is a fundamental method in machine learning classification.
@@
Many approaches to accelerate similarity queries have been proposed, most of them based on access methods. These methods use efficient indexing structures that divide the representation space into regions defined by a distance function. Access methods include a wide range of techniques, ranging from exact methods to approximate solutions, and adaptations designed for distributed computing systems.
@@ 2.2 Algorithms for similarity search
Exact similarity search methods identify the completely accurate result set for a query by using the absolute positions of objects to group elements and search for similar instances. These methods are categorised as Point Access Methods (PAM) for point-based elements, and Spatial Access Methods (SAM) for elements with spatial dimensions, such as lines, polygons or higher-dimensional polyhedra. The general acronym SAM is commonly used in the research literature to refer to both categories of indexing strategies in vector spaces [14]. These approaches are most effective with elements represented in a relatively small number of dimensions, making the Curse of Dimensionality in high-dimensional spaces one of the biggest challenges in data science today.
@@
Metric Access Methods (MAM) provide a general framework for similarity search problems within metric spaces [11, 13], exploiting specific properties like the triangle inequality to partition the metric space into subspaces, effectively filtering out portions of the dataset that cannot contain valid results for a query. On the one hand, MAM exact search can be expensive in terms of computational time and updating the index structure due to dynamic data. On the other hand, MAM indexes have serious problems in efficiently reducing the search for candidate elements in a query when dealing with high-dimensional data. There are many MAM for queries in metric spaces, some of which use clustering algorithms, such as $k$-means [19] or $k$-medoids [20], to recursively partition the entire set of elements creating a hierarchical structure (tree) of nested clusters.
Approximate indexing methods strike a balance between accuracy and speed, making them applicable to both vector and metric spaces [21]. Despite potential inaccuracies in the results, they provide a faster response by minimising distance calculations. These methods are particularly useful in contexts where exact approaches are hampered by the computational burden of exhaustive scans, especially in high-dimensional datasets or scenarios involving large amounts of data
Rather than computing the exact result set for a query, approximate similarity search algorithms use various strategies (e.g., transforming the feature space, reducing the number of pairwise distances computed) to identify a reduced set of candidate items that are likely to be highly similar. Some of the most commonly used approximate algorithms for speed up similarity search are tree-based methods [22], hashing-based methods [23] and graph-based methods [24].
Tree-based methods divide the high-dimensional space hierarchically, grouping similar objects in the same partition as leaf nodes. In addition, objects called representatives are often used to represent each partition. Thus, each node stores the representatives, the objects in the covered partition, and their distances to the representatives. When a search query is performed, the query object is first compared to the representatives of each partition, which reduces the overall search time. Popular tree-based methods include Vantage Point Tree (VP-tree) [25], M-tree[26] and Slim-Tree [27]. More recently, several methods have been proposed to optimise the structure of the disk index, such as HD-Index [28], or methods that focus on parallel or distributed processing, such as Hercules [29] and TARDIS [30], respectively.
Hashing-based methods aim to transform objects into low-dimensional representations, called hash codes [31]. The hash codes are short sequences of bits designed to preserve the similarity between the original objects, ensuring that similar objects are mapped to similar codes. These methods are typically divided into data dependent and data independent approaches. Data dependent methods, based on machine learning or deep learning algorithms, use the underlying structure and distribution of data to design hash functions, resulting in more accurate and meaningful hash codes. Some methods based on this approach are BoostMap [32], Sequential Projection Learning Hashing (SPLH) [33], K-Means Hashing (KMH) [34] and Deep Hashing Network architecture for supervised hashing (DHN) [35]. In contrast, data independent methods, such as LSH [4], use random projections to map samples in a feature space without requiring detailed knowledge of the data.
Graph-based methods exploit the inherent structure of data by representing it as a graph, where nodes correspond to data points and edges represent the similarity relationship between the connecting pair of nodes [24]. This approach facilitates the retrieval of similar objects through graph traversal algorithms, which can efficiently narrow down the search space. Several graph methods have been proposed, such as $k$-Nearest Neighbour Graph (KNNG) [36], NN-Descent [37], Hierarchical Navigable Small World Graph (HNSW) [38], Fast Approximate Nearest Neighbour Graph (FANNG) [39] and Navigating Spreading-out Graph (NSG) [40].
@@ 2.3 Using different distance functions
The dissimilarity value captures the notion of closeness between two objects, which is interpreted as a distance in a suitable space. A distance function that avoids introducing any correlation between features, such as the Minkowski distance [41] or $\ell_{p}$ norm, can quantify this value [42]. The Manhattan or city block distance ($\ell_{1}$ norm), the Euclidean distance ($\ell_{2}$ norm), and the Chebyshev distance ($\ell_{\infty}$ norm) are typical instances of this family. Many indexing structures and algorithms have commonly relied on the Euclidean distance, derived from its conventional use in two- or three-dimensional spatial applications.
@@
However, several drawbacks arise when dealing with high-dimensional problems, an effect widely known as the Curse of Dimensionality. As the number of dimensions increases, the volume of the space grows exponentially, and the density of points in the space decreases. This sparsity poses a challenge in accurately measuring distances and similarities between data elements, since proximity becomes less meaningful. Consequently, elements can be very far apart, even though they are close according to a metric such as the Euclidean distance. This effect increases the computational complexity of algorithms and operations performed in high-dimensional spaces, often resulting in slower and less efficient computations. For this reason, several authors have pointed out that, due to the Curse of Dimensionality, it is very costly to find the exact nearest neighbour in high-dimensional Euclidean spaces [43, 44].
@@
As a result, although Euclidean spaces reflect an intuitive conception of distance, it is challenging to extrapolate this natural assumption with high-dimensional data. Moreover, traditional clustering algorithms, such as $k$-means, often specify the Euclidean distance as the preferred metric without considering other possibilities. This choice brings in a self-imposed limitation that may produce results far from the best [45, 46, 9].
@@
In recent years, several surveys [47, 48, 49] have addressed the evaluation of alternative metrics for similarity searches across spaces of diverse topology, a burgeoning topic offering insights into the most suitable similarity measures for different problems and contributing to the ongoing development of new methods for similarity search.
@@
Therefore, we aim to introduce GDASC, an algorithm that enables the integration of arbitrary distances through the use of data partitioning algorithms that generate Voronoi regions and assign a prototype for each of these regions. Specifically, the utilisation of several clustering algorithms, such as $k$-medoids, allows for the incorporation of multiple $\ell_{p}$ norms or even other metrics for specific contexts, facilitating parametrisation with different distances and the execution of validation experiments. Our findings demonstrate the feasibility of employing diverse distances beyond the Euclidean metric, highlighting the potential for achieving superior results in data indexing, contingent upon the specific characteristics of the dataset.